The air pollutant NO is produced in automobile engines from the high-temperature reaction N2(g)+O2(g)⇌2NO(g);Kc=1.7×10−3 at 2300 K.

If the initial concentrations of N2 and O2 at 2300 K are both 1.58 M, what are the concentrations of NO when the reaction mixture reaches equilibrium?

Concentrations of NO= UNITS

If the initial concentrations of N2 and O2 at 2300 K are both 1.58 M, what are the concentrations of N2 when the reaction mixture reaches equilibrium?

Concentrations of O2=

  If the initial concentrations of N2 and O2 at 2300 K are both 1.58 M, what are the concentrations of O2 when the reaction mixture reaches equilibrium?

Concentrations of O2=
Answer #1

NO C C7) ー ( l-S8-x) (1.58-x) ㄐㄨ 2 7996 [nJ = [cz〕 = l-SS-o.oo 2126 = 1.573 m